A recent study of 300 patients found that of 100 alcoholic patients, 87 had elevated cholesterol levels, and of 200 nonalcoholic patients, 43 had elevated cholesterol levels. If a patient is selected at random, find the probability that the patient is the following.

An alcoholic with elevated cholesterol level

A nonalcoholic

A nonalcoholic with non-elevated cholesterol level


1
Expert's answer
2022-01-24T16:35:09-0500
"P(A)=\dfrac{100}{300}=\dfrac{1}{3}"

"P(\bar{A})=1-P(A)=\dfrac{2}{3}"

"P(Ch|A)=\dfrac{87}{100}=0.87"

"P(Ch|\bar{A})=\dfrac{43}{200}=0.215"



a)


"P(A\cap Ch)=P(A)P(Ch|A)=\dfrac{1}{3}(0.87)=0.29"

b)


"P(\bar{A})=1-P(A)=\dfrac{2}{3}"

c)


"P(\bar{A}\cap \overline{Ch})=P(\bar{A})P(\overline{Ch}|\bar{A})=\dfrac{2}{3}(1-0.215)"

"=\dfrac{157}{300}\approx0.5233"